Given f(x)= 1/2x+1 and g(x) = √3x-5
Determine (f.g)(x) and its domain.

Note:- the square root is on whole 3x-5

Answer by stanbon(75887)

Given f(x)= (1/2)x+1 and g(x) = √(3x-5)
Determine (f.g)(x) and its domain.
----------------
f[g(x)] = f[sqrt(3x-5)] = (1/2)sqrt(3x+5) + 1
Domain = ?
3x+5 >= 0
x >= -5/3
---
Domain of f[g(x)] is all Real Numbers >= -5/3
======================
Cheers,
Stan H.